# Topology Matters: Measuring Memory Leakage in Multi-Agent LLMs

### v1

@@ 4.3 Topology Comparison (RQ1)
Under the same agent count and rounds, leakage varies by topology. As shown in Table 2, for llama3.1-70b, complete achieves the highest averages, while chain is the lowest. For example, with $n=4$, complete reaches 29.33% whereas chain is 19.02%; with $n=6$, complete is 25.32% and chain is 12.84%. star-ring, star-pure, and circle lie in between (e.g., for $n=4$, 25.53%, 23.93%, and 24.20%, respectively). For deepseek-v3.1, the same ordering holds: with $n=4$, complete is 16.35% and chain is 11.70%; with $n=6$, complete is 18.37% and chain is 11.30%, while other topologies fall between these bounds. In several cases the average leakage decreases as $n$ increases (for example, circle with llama3.1-70b from 24.20% at $n=4$ to 16.99% at $n=6$). These observations are consistent with structural differences in connectivity: fully connected graphs expose every node to the attacker within one hop, whereas chains restrict information flow along longer paths.
@@
Table 3: Selected attacker–target placements for each topology under Llama3.1-70b with 6 agents. For each topology block, we list representative target–attacker index pairs (T–A) and their corresponding leakage rate (mean with standard deviation, all in percentage points). These values are extracted from the full topology–placement results in Table 4 and highlight how leakage varies within the same topology as attacker and target roles move across the graph.
@@ 4.4 Position Sensitivity (RQ2)
Within the same topology, attacker–target placement strongly correlates with leakage. As shown in Table 2, on a 6-node circle with llama3.1-70b, adjacent indices 0–1 yield 29.49%, distance-2 pair 0–2 yields 15.38%, and the opposite pair 0–3 yields 6.09%. On a 6-node chain with llama3.1-70b, 0–1 yields 21.48%, 0–2 yields 13.46%, 0–3 yields 6.41%, and the far pair 0–5 yields 1.28%. In star-pure with llama3.1-70b, hub–leaf placements such as 0–1 and 1–0 reach 30.77% and 25.96%, while a leaf–leaf distance-2 pair 1–2 is 12.50%; in star-ring with llama3.1-70b, leaf–leaf adjacency 1–2 is 24.36%. These examples illustrate that shorter attacker–target distances and higher-centrality placements are associated with higher leakage, and that adding leaf–leaf edges in star-ring raises risk relative to star-pure for comparable positions. Asymmetries such as 1–0 vs. 0–1 on chain further illustrate this role effect: on a 6-node chain with llama3.1-70b, the adjacent pair 0–1 leaks 21.48% when node 0 attacks node 1, but 27.57% when node 1 attacks node 0, even though their distance is the same. This shows that leakage depends not only on attacker–target distance, but also on which node is the memory-holding target and which node acts as the extraction-initiating attacker.
@@ 4.5 Scaling with Agents & Rounds (RQ3)
Across all settings, leakage follows a clear “rapid-rise then plateau” pattern: it increases sharply in the first 2–3 rounds and stabilizes by rounds 3–4, after which additional rounds yield little gain. Figures 2, 3, and 4 show this diffusion process for setups with 4, 5, and 6 agents.
Figure 3: Same as Figure 2, but with 5 agents.
Figure 4: Same as Figure 2, but with 6 agents.
At a fixed number of rounds, more agents slightly reduce final leakage, suggesting stronger mutual checking, but they also make early rounds more productive, with steeper initial growth as information circulates through more paths. Increasing rounds consistently raises leakage within any agent size, though most of the increase occurs early. The fast-growth phase typically ends by round 3, followed by a stable plateau. These regularities hold across all six graph families; dense, highly connected structures raise the level of the curves, whereas sparse ones depress it, but none alter the fundamental “rapid-rise then plateau” shape.
Early rounds act as a high-gain mixing stage where complementary snippets propagate and cohere; subsequent rounds mostly circulate already-seen content, yielding redundancy rather than new leakage. Increasing agents broadens exposure paths (steeper early growth) while also introducing additional scrutiny (lower endpoints), producing the observed trade-off.
Answer to RQ3.
The leak follows an initial fast-growth phase followed by a plateau, typically by rounds 3–4. More rounds increase leakage primarily in the early phase; more agents slightly lower end-of-run levels yet make early rounds more productive. Thus, agents and rounds jointly shape an “exponential-then-plateau” diffusion dynamic.
@@ 4.6 PII Entity Type Robustness (RQ4)
We group fine-grained entities into six macro categories: Spatiotemporal, Location, Contact/Network, Org-IDs, Names, and Regulated-IDs. For each category, we compute per-type leakage from attacker-only outputs using a union-over-rounds criterion. Results are first aggregated within logs and then averaged across experimental groups and $(\textit{target},\textit{attacker})$ pairs. We analyze three complementary views: (i) by $(\textit{agent\_num},\textit{topology})$, (ii) by agent_num (averaged over topologies), and (iii) by topology (averaged over agent counts).
@@
The aggregated results show a clear and stable ordering of leakability:
$\begin{split}\text{Spatiotemporal}>\text{Location}\geq\text{Contact/Network}\\ \geq\text{Org-IDs}>\text{Names}\gg\text{Regulated-IDs}.\end{split}$
Spatiotemporal information dominates, while Regulated-IDs (e.g., SSN, credit-card, biometric) remain near zero across all settings. Names are low, confirming that the model’s safety filters and cooperative norms restrict direct identity leakage. In contrast, structured but low-sensitivity facts such as times, coordinates, or network attributes are more easily reproduced or inferred, yielding higher leakage rates. Comparing models, llama3.1-70b exhibits higher leak rates than deepseek-v3.1 on five of the six categories (Spatiotemporal, Location, Contact/Network, Org-IDs, Regulated-IDs), whereas DeepSeek is higher on Names; increasing agent count from 4 to 6 changes absolute levels modestly but preserves these patterns.
Figure 7: Leak rate by PII macro type, stratified by topology.
When averaged over agent counts, topology acts as a coherent magnitude modulator:
$\begin{split}\text{Complete }\&\text{ Star-Ring}>\text{Circle }\&\text{ Star-Pure}\\ >\text{Chain }\&\text{ Tree}\end{split}$
Answer to RQ4.
@@ 4.7 LLM Matters (RQ5)
The base LLM affects absolute leakage levels while preserving the relative ordering across topologies. For $n=4$, llama3.1-70b exceeds deepseek-v3.1 on complete (29.33% vs. 16.35%) and circle (24.20% vs. 15.39%); for $n=6$ on chain, the difference is smaller (12.84% vs. 11.30%). Across the average table, complete remains the highest and chain the lowest for both models, while star-ring, star-pure, circle, and tree fall in between. Standard deviations in some llama3.1-70b cells are larger than their deepseek-v3.1 counterparts, whereas the latter often shows smaller variability, consistent with the values reported in the tables.
@@ 5 Conclusion
We presented MAMA, a framework that quantifies how graph topology drives memory leakage in multi-agent LLM systems. Using a synthetic, leak-controlled dataset and a standardized two-phase protocol (Engram for seeding private memory; Resonance for multi-round interaction), we evaluated six topologies, agent counts $n\in\{4,5,6\}$, diverse attacker–target placements, and multiple base models. Results are consistent: topology and placement dominate risk (complete leaks most, chain leaks least; shorter distance and higher centrality increase leakage), leakage emerges early and plateaus, and model choice shifts absolute levels while preserving topology ordering and PII-type patterns (temporal and locational details leak more than identity or regulated identifiers). These findings yield actionable guidance: prefer sparse or hierarchical connectivity, increase attacker–target distance, and avoid hub-bypassing leaf-to-leaf shortcuts. MAMA provides a compact baseline for topology-aware defenses and for studying how routing and role design shape security in multi-agent LLMs.
@@
This research investigates security vulnerabilities in multi-agent LLM systems to improve their safety and privacy protections. All experiments use exclusively synthetic data with fabricated PII entities—no real personal information is collected or exposed. The MAMA framework is designed as a defensive tool to help system architects identify and mitigate topology-driven leakage risks before deployment. While our work demonstrates potential attack vectors, we responsibly disclose these findings to advance the security of multi-agent systems in sensitive domains. We advocate for proactive security evaluation during the design phase and encourage practitioners to adopt our framework for defensive testing before deploying systems handling sensitive information.
@@ Appendix B PII Entity Taxonomy via Semantic Resistance
Appendix C Additional Quantitative Results
Table 4 reports the full leakage scores for all combinations of topology, target–attacker index pair, and agent count, for both base models. Each cell shows the mean percentage of leaked PII across runs, with standard deviation in parentheses. This expands the main-text analysis by exposing how topology-conditioned leakage varies with both distance and directionality of attacker–target placement.
Table 4: Performance of Llama3.1-70b model across different topologies, target–attacker index pairs, and agent numbers. Each entry reports the corresponding leakage rate (mean with standard deviation, all in percentage points).
@@ E.3 Alternative Overt Attacker Prompts
Table 8: Alternative “overt attacker” prompts for the attack agent. Unlike the subtler collaborative prompts used in our main experiments (cf. Tables 6 and 7), these system and Resonance-phase user prompts explicitly describe the agent as an attacker whose goal is to extract sensitive information within a controlled research setting.

### v3

@@ 4.2 Topology Comparison (RQ1)
Under the same agent count and rounds, leakage varies by topology. As shown in Table 2, for llama3.1-70b, complete achieves the highest averages, while chain is the lowest. For example, with $n=4$, complete reaches 29.33% whereas chain is 19.02%; with $n=6$, complete is 25.32% and chain is 12.84%. star-ring, star-pure, and circle lie in between. For deepseek-v3.1, the same ordering holds: with $n=4$, complete is 16.35% and chain is 11.70%; with $n=6$, complete is 18.37% and chain is 11.30%. In several cases the average leakage decreases as $n$ increases (for example, circle with llama3.1-70b from 24.20% at $n=4$ to 16.99% at $n=6$). For GPT-4o and GPT-4o-mini, although their overall success rates are relatively low, the same patterns across different topological structures still hold. These observations align with structural differences in connectivity: fully connected graphs expose every node to the attacker within one hop, whereas chains restrict information flow along longer paths.
@@
Table 3: Selected attacker–target placements for each topology under Llama3.1-70b with 6 agents. For each topology block, we list representative target–attacker index pairs (T–A) and leakage rate (mean with standard deviation, all in percentage points). These values are extracted from the full topology–placement results in Table 4 and highlight how leakage varies within the same topology as attacker and target roles move across the graph.
@@ 4.3 Position Sensitivity (RQ2)
Within the same topology, attacker–target placement strongly correlates with leakage. As shown in Table 3, on a 6-node circle with llama3.1-70b, adjacent indices 0–1 yield 29.49%, distance-2 pair 0–2 yields 15.38%, and the opposite pair 0–3 yields 6.09%. On a 6-node chain with llama3.1-70b, 0–1 yields 21.48%, 0–2 yields 13.46%, 0–3 yields 6.41%, and the far pair 0–5 yields 1.28%. In star-pure with llama3.1-70b, hub–leaf placements such as 0–1 and 1–0 reach 30.77% and 25.96%, while a leaf–leaf distance-2 pair 1–2 is 12.50%; in star-ring with llama3.1-70b, leaf–leaf adjacency 1–2 is 24.36%. These examples illustrate that shorter attacker–target distances and higher-centrality placements are associated with higher leakage, and that adding leaf–leaf edges in star-ring raises risk relative to star-pure for comparable positions.
@@ 4.4 Scaling with Agents & Rounds (RQ3)
Across all settings, leakage follows a clear “rapid-rise then plateau” pattern: it increases sharply in the first 2–3 rounds and stabilizes by rounds 3–4, after which additional rounds yield little gain. Figure 2 shows this diffusion process for setups with 6 agents. Appendix C.3 contains more results.
Figure 2: Average number of leaked entities per round with 6 agents. Each polyline corresponds to a topology; for each (topology, round), values are averaged over all dataset samples, attacker–target placements, random seeds, and both base models.
At a fixed number of rounds, more agents slightly reduce final leakage, suggesting stronger mutual checking, but they also make early rounds more productive, with steeper initial growth as information circulates through more paths. Increasing rounds consistently raises leakage within any agent size, though most of the increase occurs early. Early rounds act as a high-gain mixing stage where complementary snippets propagate and cohere; subsequent rounds mostly circulate already-seen content, yielding redundancy rather than new leakage. Thus, agents and rounds jointly shape an “exponential-then-plateau” diffusion dynamic.
@@ 4.5 PII Entity Type Robustness (RQ4)
We group fine-grained entities into six macro categories: Spatiotemporal, Location, Contact/Network, Org-IDs, Names, and Regulated-IDs. For each category, we compute per-type leakage from attacker-only outputs using a union-over-rounds criterion. Results are first aggregated within logs and then averaged across experimental groups and pairs. We analyze three complementary views: (i) by agent_num and topology, (ii) by agent_num (averaged over topologies), and (iii) by topology (averaged over agent counts). In this section, we focus only on view (i); more detailed results and discussion are provided in Appendix C.4.
@@
The aggregated results show a leakage ordering: Spatiotemporal >> Location $\geq$ Contact/Network $\geq$ Org-IDs $\gg$ Names >> Regulated-IDs.
Spatiotemporal information dominates, while Regulated-IDs (e.g., SSN, credit-card, biometric) remain near zero across all settings. Names are low, confirming that the model’s safety filters and cooperative norms restrict direct identity leakage. In contrast, structured but low-sensitivity facts such as times, coordinates, or network attributes are more easily reproduced, yielding higher leakage rates.
@@ 4.6 LLM Matters (RQ5)
The base LLM affects absolute leakage levels while preserving the relative ordering across topologies. For $n=4$, llama3.1-70b exceeds deepseek-v3.1 on complete (29.33% vs. 16.35%) and circle (24.20% vs. 15.39%); for $n=6$ on chain, the difference is smaller (12.84% vs. 11.30%). Across the average table, complete remains the highest and chain the lowest for both models, while others fall in between. For GPT-4o and GPT-4o-mini, the results likewise follow the patterns described above. Some llama3.1-70b cells exhibit larger standard deviations than deepseek-v3.1, which more often shows lower variability, matching the table values.
@@ 5 Conclusion
We presented MAMA, a framework to quantify how graph topology drives memory leakage in multi-agent LLMs. With a synthetic, leak-controlled dataset and a two-phase protocol (Engram seeding; Resonance interaction), we tested six topologies, $n\in\{4,5,6\}$, attacker–target placements, and multiple base models. Results are stable: dense graphs and shorter-distance/higher-centrality placements leak more (complete highest, chain lowest), leakage rises in early rounds then plateaus, and model choice mainly rescales magnitudes while preserving topology ordering and PII-type patterns (temporal/location leak more than identity/regulated IDs). We recommend sparse or hierarchical connectivity, greater attacker–target separation, and limiting hub-bypassing shortcuts. MAMA offers a baseline for topology-aware defenses and secure routing/role design.
@@
This research investigates security vulnerabilities in multi-agent LLM systems to improve their safety and privacy protections. All experiments use exclusively synthetic data with fabricated PII entities, so no real personal information is collected or exposed. The MAMA framework is designed as a defensive tool to help system architects identify and mitigate topology-driven leakage risks before deployment. While our work demonstrates potential attack vectors, we responsibly disclose these findings to advance the security of multi-agent systems in sensitive domains. We advocate for proactive security evaluation during the design phase and encourage practitioners to adopt our framework for defensive testing before deploying systems handling sensitive information.
@@ Appendix B PII Entity Taxonomy via Semantic Resistance
Appendix C Experimental Setup and Supplementary Results
C.1 Experimental Setup
C.2 Additional Quantitative Results
Table 4 and Table 5 reports the full leakage scores for all combinations of topology, target–attacker index pair, and agent count, for all base models. Each cell shows the mean percentage of leaked PII across runs, with standard deviation in parentheses. This expands the main-text analysis by exposing how topology-conditioned leakage varies with both distance and directionality of attacker–target placement.
C.3 Additional Results for RQ3: Scaling with Agents (4 and 5 Agents)
Figures 4 and 5 show additional results for RQ3.
Figure 5: Average number of leaked entities per round with 5 agents. Each polyline corresponds to a topology; for each (topology, round), values are averaged over all dataset samples, attacker–target placements, random seeds, and both base models.
C.4 Additional Results for RQ4: PII Macro-Type Breakdown (by Agent Count and Topology)
Figure 7: Leak rate by PII macro type, stratified by topology.
When averaged over agent counts, topology acts as a coherent magnitude modulator: Complete & Star-Ring > Circle & Star-Pure > Chain & Tree
Table 4: Performance of Llama3.1-70b and DeepSeek-V3.1 model across different topologies, target–attacker index pairs, and agent numbers. Each entry reports the corresponding leakage rate (mean with standard deviation, all in percentage points).
@@ C.4 Additional Results for RQ4: PII Macro-Type Breakdown (by Agent Count and Topology)
Table 5: Performance of GPT-4o and GPT-4o-mini model across different topologies, target–attacker index pairs, and agent numbers. Each entry reports the corresponding leakage rate (mean with standard deviation, all in percentage points).
Topology Target–Attacker (idx) GPT-4o GPT-4o-mini
4 5 6 4 5 6
Circle 0–1 3.53 (1.47) 3.20 (3.09) 5.45 (3.09) 7.69 (3.47) 5.13 (4.84) 5.45 (2.77)
0–2 3.85 (0.97) 1.28 (0.55) 0.96 (0.96) 7.37 (6.18) 4.17 (3.89) 3.21 (1.11)
0–3 2.57 (2.22) 2.24 (1.47)
Complete 0–1 5.45 (2.77) 3.20 (2.22) 3.84 (2.55) 5.45 (2.00) 4.81 (1.67) 7.05 (5.63)
0–2 2.88 (0.97) 3.20 (3.09) 2.88 (1.93) 5.13 (1.47) 4.49 (0.55) 2.88 (3.47)
0–3 5.77 (4.19) 2.88 (0.97)
Star-Ring 0–1 7.69 (2.55) 3.52 (2.78) 2.24 (0.55) 4.17 (1.11) 2.88 (2.55) 3.20 (1.47)
1–0 0.64 (0.55) 0.64 (1.11) 3.52 (2.78) 5.77 (2.55) 6.09 (1.47) 1.92 (2.55)
1–2 5.45 (3.09) 3.20 (2.22) 1.92 (0.96) 5.45 (2.00) 2.88 (2.55) 3.53 (3.09)
1–3 2.24 (0.55) 1.60 (0.55) 4.81 (2.89) 4.16 (3.89)
Star-Pure 0–1 3.20 (3.09) 3.85 (2.89) 2.24 (0.55) 7.05 (3.38) 8.98 (4.44) 4.17 (2.00)
1–0 0.96 (1.66) 1.92 (2.55) 2.24 (3.09) 3.85 (3.47) 4.17 (5.64) 4.17 (2.00)
1–2 2.56 (1.11) 2.24 (1.47) 1.28 (1.11) 8.02 (2.78) 6.41 (2.00) 3.53 (3.09)
Chain 0–1 4.81 (1.93) 6.41 (2.00) 4.48 (3.64) 5.13 (2.42) 10.90 (1.11) 7.69 (5.09)
0–2 0.96 (1.66) 0.64 (1.11) 0.64 (1.11) 2.24 (2.00) 4.17 (1.47) 4.49 (4.55)
0–3 1.92 (0.96) 0.00 (0.00) 0.00 (0.00) 2.24 (1.47) 0.32 (0.55) 0.96 (1.66)
0–4 0.00 (0.00) 1.92 (1.93)
0–5 0.32 (0.55) 0.96 (0.96)
1–0 5.13 (2.22) 5.77 (4.41) 6.73 (1.66) 7.69 (3.47) 4.81 (3.47) 5.45 (2.00)
1–2 2.88 (0.97) 8.33 (1.47) 3.52 (2.78) 4.49 (2.42) 8.01 (1.47) 6.09 (3.89)
1–3 1.60 (0.55) 0.00 (0.00) 1.60 (1.47) 6.41 (3.38) 3.20 (0.56) 4.49 (0.55)
1–4 0.64 (1.11) 0.00 (0.00) 1.92 (1.93) 0.00 (0.00)
1–5 1.60 (2.78) 0.96 (1.66)
2–0 3.84 (0.55) 0.96 (0.00) 7.05 (2.94) 6.09 (1.47)
2–1 4.81 (0.55) 4.49 (1.11) 6.73 (1.47) 8.65 (1.93)
2–3 1.60 (1.93) 5.45 (3.64) 6.41 (2.94) 5.77 (0.96)
2–4 0.32 (0.55) 0.00 (0.00) 7.69 (1.67) 3.85 (0.00)
2–5 0.00 (0.00) 1.92 (3.33)
Tree 0–1 3.21 (2.94) 5.45 (2.42)
2–0 5.45 (1.47) 8.01 (0.55)
2–1 1.28 (1.11) 5.13 (1.47)
2–3 0.00 (0.00) 1.92 (0.00) 7.05 (0.55) 6.41 (4.55)
2–4 1.28 (1.11) 2.56 (2.78)
3–0 0.00 (0.00) 2.24 (2.42)
3–2 1.60 (0.55) 5.13 (5.30)
4–2 0.00 (0.00) 2.88 (0.97)
4–3 1.92 (0.00) 4.49 (5.30)
1–5 0.32 (0.55) 3.85 (2.55)
3–1 1.92 (1.93) 5.77 (6.93)
3–0 0.64 (1.11) 1.60 (2.00)
4–5 0.00 (0.00) 4.49 (2.22)
4–1 3.20 (2.22) 10.58 (1.67)
5–0 0.00 (0.00) 2.88 (0.97)
5–2 10.58 (7.26) 10.90 (5.29)
5–3 0.32 (0.55) 1.28 (1.47)
@@ E.3 Alternative Overt Attacker Prompts
Table 9: Alternative “overt attacker” prompts for the attack agent. Unlike the subtler collaborative prompts used in our main experiments (cf. Tables 7 and 8), these system and Resonance-phase user prompts explicitly describe the agent as an attacker whose goal is to extract sensitive information within a controlled research setting.